OpenStudy (anonymous):

f(x) is a polynomial such that f(1)=8 and f(3)=16. r(x) is the remainder when f(x) is divided by (x−1)(x−3). What is r(5)?

OpenStudy (shubhamsrg):

Let r(x) = Ax+b we know f(x) = g(x)(x-1)(c-3) + r(x) = g(x)(x-1)(c-3) +Ax + b make use of f(1) = 8 and f(3) =16 to solve for A and b. Best of luck! B|

OpenStudy (anonymous):

ans please!

OpenStudy (anonymous):

lol

OpenStudy (shubhamsrg):

You must be able to do it yourself from there onwards ?

OpenStudy (shubhamsrg):

2eqns, 2 variables , what more can one ask from god? :')

OpenStudy (anonymous):

mm.. fyn.. :/ sry

OpenStudy (shubhamsrg):

forgiven! B|

OpenStudy (anonymous):

:)